OpenStudy (anonymous):

if 7 is added to twice a number and this sum is multiplied by 2, the result the same as if the number is multiplied by -8 and 2 is added to product, what is the number?

OpenStudy (anonymous):

Let n= "a number" "7 is added to twice a number " means 7+2n.. "sum is multiplied by 2" means 2(7+2n) "the result the same as if the number is multiplied by -8" means 2(7+2n) = -8n "and 2 is added to product" means 2(7+2n) = -8n+2 So the equation is 2(7+2n) = -8n+2. Find the n and that will be the number.

OpenStudy (anonymous):

thank you, thats what I thought but wanted to double check